A body falls freely under gravity.The distance travelled by it in the last second of its journey equals the distance travelled by it in the first three seconds. the total time off all is(a) 5s (b) 8s (c) 12s (d) 15s

Grade:12th pass

1 Answers

Piyush Kumar Behera
7 years ago
The answer is 5 seconds..
Distance travelled in first three seconds=s=1/2at2=1/2 * 10 *32=45m
Distance travelled in nth second=u + g(2n-1)/2
as the magnitude of last second is the magnitude of the total time taken.
Distance travelled in nth second=Distance travelled in first three seconds=45m
=>u + g(2n-1)/2=45
as u=0
=> g(2n-1)/2=45
=> n=5
so the last second is 5th second so the total time taken is 5 seconds.
I Hope you understood the solution!!Please upvote the solution to encourage me..